Question) Jase Manufacturing Co.'s static budget at 7,800 units of production includes $23,400 for direct labor and $2,340 for electric power. Total fixed costs are $41,200. At 10,500 units of production, a flexible budget would show?

A)variable and fixed costs totaling $66,940

B)variable costs of $34,650 and $41,200 of fixed costs

C) variable costs of $34,650 and $55,462 of fixed costs

D)variable costs of $25,740 and $41,200 of fixed costs
Answer #1

Total variable cost=(23400+2340)=$25740

Hence variable cost/unit=(25740/7800)=$3.3

Hence total variable cost at 10500 units=(3.3*10500)=$34650

Hence the correct option is B.

NOTE:Total fixed costs and variable cost/unit do not change with change in units.